What is the duration of a 10-year bond, selling for $877.11, with an 8% annual coupon? a. 7.62 b.7.33 c.7.04 d. 6.95
Answer #1

Par Value of Bond = $1,000

Annual Coupon Rate = 8%

Hence, coupon amount = $80

Let us depict the yearly cash flows given by this bond :-

А B в C D Years Cash Flow Hoc vouw 1080

Bond is currently trading at $877.11.

Let us use "Goal Seek" function to find out the discount rate for arriving at the current price of bond

Using Goal seek we find that discount rate =10%

Attaching the calculations:-

А в C D Discount Rate 10% Years 0.68 Cova un AwN Cash Flow PV Factor PV of Cash Flows 800.91 72.73 0.83 66.12 0.75 60.11 54.6

A B Discount Rate 0.1 80 2 80 Years Cash Flow PV Factor =1/(1+$D$2) B5 =1/(1+$D$2) B6 =1/(1+$D$2)^B7 =1/(1+$D$2)^B8 =1/(1+$D$

Now, let us calculate the Present Value of Weighted Cash Flow, by multiplying the "PV of Cash Flows" column with Years column.

Attaching the calculations:-

A B C D E Discount Rate 10% Years Cash Flow PV Factor PV of Cash Flows PV of Weighted Cash Flow 800.91 72.73 72.73 80 0.83 66

A B _c C Discount Rate 0.1 80 Years Cash Flow PV Factor 80 =1/(1+$D$2)^B5 =1/(1+$D$2)^B6 =1/(1+$D$2)^B7 =1/(1+$D$2)^B8 =1/(1+

To arrive at the Duration we have to divide the sum of "PV of Weighted Cash Flow" by the current market price.

Hence, Duration = (6,178.31 / 877.11) = 7.04 years

Answer is Option C